# ToMoE: Converting Dense Large Language Models to Mixture-of-Experts through Dynamic Structural Pruning

### v1

@@ Abstract
Large Language Models (LLMs) have demonstrated remarkable abilities in tackling a wide range of complex tasks. However, their huge computational and memory costs raise significant challenges in deploying these models on resource-constrained devices or efficiently serving them. Prior approaches have attempted to alleviate these problems by permanently removing less important model structures, yet these methods often result in substantial performance degradation due to the permanent deletion of model parameters. In this work, we tried to mitigate this issue by reducing the number of active parameters without permanently removing them. Specifically, we introduce a differentiable dynamic pruning method that pushes dense models to maintain a fixed number of active parameters by converting their MLP layers into a Mixture of Experts (MoE) architecture. Our method, even without fine-tuning, consistently outperforms previous structural pruning techniques across diverse model families, including Phi-2, LLaMA-2, LLaMA-3, and Qwen-2.5.
Machine Learning, ICML
@@ 1 Introduction
Large Language Models (LLMs) have emerged as powerful tools in artificial intelligence, exhibiting remarkable capacity in natural language understanding and generation. By leveraging advanced neural architectures (Vaswani et al., 2017), and pretraining on vast corpora of text, these models have demonstrated a remarkable ability to perform diverse tasks, ranging from question answering and machine translation to content creation and scientific discovery (Brown, 2020; Kenton & Toutanova, 2019; Raffel et al., 2020). Examples of popular commercial or non-commercial LLMs include ChatGPT (OpenAI, 2022), Claude (Anthropic, 2023), and LLaMA (Touvron et al., 2023a). The core strength of LLMs lies in their ability to capture complicated relationships within the context, enabling them to generalize across tasks with in-context learning (Dong et al., 2022) or minimal fine-tuning (Radford et al., 2019; Kaplan et al., 2020).
Figure 1: (a): The original LLM uses all parameters to process all the input text. (b): The static pruning for LLMs permanently removes model parameters, and the rest of the parameters are used to process all the input text. Our approach (c): LLMs with dynamic pruning use different sub-networks to process different tokens. We incorporate MoE to achieve a fixed expected budget in inference.
@@
Although LLMs demonstrate remarkable potential, their huge model size often limits their usability on devices with limited resources. As a result, considerable efforts (Ma et al., 2023; Ashkboos et al., 2024; Frantar et al., 2022) are focused on minimizing the computational and memory costs of these models. Structural pruning (Ma et al., 2023) has emerged as a promising solution to this challenge because, unlike unstructured pruning, it achieves compression without the need for specialized implementations. However, the problem with structural pruning methods is that they will substantially reduce the model capacity resulting in an obvious performance gap compared to the dense model. The fine-tuning cost to even partially recover this gap is tremendous. To achieve a better trade-off between the number of parameters and performance, sparse Mixture of Experts (MoE) models (Shazeer et al., 2017; Lepikhin et al., 2021b) are designed to activate only a subset of the model’s parameters, corresponding to the selected experts. Recently proposed MoE models, such as DeepseekMoE (Dai et al., 2024), demonstrated that they can match the performance of dense models with a similar total parameter count while using a small number of active parameters. Following this lead, we propose that transforming dense models into MoE models could offer a promising approach to bridging the performance gap left by structural pruning methods. Unlike prior efforts to construct MoE models from dense models (Zhu et al., 2024), our findings reveal that MoEs inherently exist within dense models and can be uncovered without updating model weights (continue pretraining). Specifically, we show that these experts can be identified through dynamic structural pruning. These results represent a novel contribution that has not been demonstrated in previous studies.
@@
The core idea of MoE models is conditional computation, where experts are dynamically selected based on input tokens. This concept aligns closely with dynamic pruning methods (Gao et al., 2019), which make pruning decisions given input features. Leveraging this connection, we propose to construct MoE models from dense models by using dynamic structural pruning. Specifically, for Multi-Head self-Attention (MHA) layers, we apply top-K routing and static pruning for compression, while for MLP layers, we transform them into MoE layers using top-1 expert routing. The routing mechanism learned for dynamic structural pruning can be directly applied to serve as the routing module for MoE layers. With differentiable discrete operations, the MoE conversion process can be formulated as a differentiable dynamic pruning problem. With this formulation, we can efficiently convert a dense model to an MoE model at a cost similar to or lower than regular structural pruning methods. The comparison between our method, static pruning, and the original LLM is shown in Fig. 1. The contributions of this work can be summarized as follows:
@@
  • Joint Optimized Training for Routing and Experts: The proposed method involves jointly optimizing routing modules and expert configurations by solving a regularized optimization problem. Our approach leverages differentiable operations to enable efficient and flexible MoE constructions.
  • Consistent Performance Improvements Across Tasks and Models: Our method consistently outperforms state-of-the-art structural pruning techniques on various tasks even without fine-tuning the model weights. This performance improvement is demonstrated across widely used public models such as Phi-2, LLaMA-2, LLaMA-3, and Qwen-2.5.
@@ 2 Related Works
Pruning: Structural pruning (Li et al., 2017; Kurtic et al., 2022; Ma et al., 2023) is an attractive technique for real-world deployments since it removes redundant parameters to reduce model size without requiring specialized implementations. Structural pruning methods fall into two main categories: static pruning (Anwar et al., 2017; Molchanov et al., 2019; Fang et al., 2023) and dynamic pruning (Gao et al., 2019; Chen et al., 2020; Anagnostidis et al., 2023; Dong et al., 2024a). Static pruning removes parameters based on input-agnostic importance metrics. For example, LLM-Pruner (Ma et al., 2023) eliminates non-essential coupled structures using gradient-based criteria. The problem with structural pruning is that it often creates a noticeable performance gap relative to dense models (Ma et al., 2023; Ashkboos et al., 2024). In contrast, dynamic pruning removes weights based on input-dependant metrics. Early attempts for dynamic pruning (Gao et al., 2019; Chen et al., 2020) focus on Convolutional Neural Networks, where channels are selectively activated for input samples. Recent works, such as D-LLM (Wang et al., 2024), incorporate the concept of conditional computation into LLMs by selectively skipping layers based on input tokens. The problem with dynamic pruning methods is that they do not have a fixed budget given different inputs, which creates problems when serving LLMs in a mini-batch setting or in the prefilling stage. Our method, on the other hand, converts the dense LLM to a sparse MoE model with fix per-token budget.
@@
MoE: Sparse Mixture-of-Experts (MoE) models improve upon pure structural pruning by maintaining or even enhancing model capacity without a proportional increase in computational cost. For instance, Sparsely-Gated MoE (Shazeer et al., 2017) employs a trainable gating network to select a small subset of experts for each input, enabling the model to scale to thousands of experts efficiently (Lepikhin et al., 2021b). More recent methods like DeepSeekMoE (Dai et al., 2024) further address expert specialization, matching dense-model performance with a similar number of activated parameters. LLAMA-MoE (Zhu et al., 2024) takes a pre-trained dense LLM, splits large feed-forward layers into multiple experts, and then dynamically routes tokens to these experts via a gating network with significant additional training. Unlike LLaMA-MoE, which treats expert construction and learning routing gates (with 200B training tokens) as distinct processes, our method integrates expert construction directly into the pruning stage, eliminating the need for separate training steps.
@@ 3 ToMoE
Most recent LLMs, like GPT (Radford et al., 2018), LLaMA (Touvron et al., 2023a), etc., adapt decoder-only architectures and thus our method focuses on decoder-only architectures. A typical decoder block consists of Multi-Head Attention (MHA) and Multi-Layer Perceptron (MLP) layers. For clarity, we let $T$, $d$, $d_{\text{mid}}$, and $H$ denote the sequence length, hidden dimension, MLP middle layer dimension, and the number of attention heads, respectively.
To reduce the computational costs of the decoder-only architecture, we propose to convert the original model into MoE models. For MHA layers, we utilize top-k routing and static pruning along the head dimension $\frac{d}{H}$. Top-k routing and static pruning for MHA layers ensure that, during prefilling, all tokens maintain the same head dimension, enabling parallel processing. For MLP layers, our approach transforms them into MoE layers along the MLP middle dimension $d_{\text{mid}}$ and employs top-1 routing. A key distinction between our method and previous dynamic pruning approaches is that the converted model maintains consistent computational costs for all inputs. This property could be crucial for efficient model serving.
@@ 3.1 Expert Embeddings
Inspired by the recent success of using hypernetworks (Ha et al., 2016; Ganjdanesh et al., 2024; Gao et al., 2024) to generate pruning decisions, we adopt a hypernetwork to generate expert embeddings:
[…]
where the $z$ is the input to the hypernetwork drawn from a random distribution, and $\mathbf{E}_{\text{all}}=[\mathbf{E}_{1},\cdots,\mathbf{E}_{l},\cdots,\mathbf{E}_{L}]$ contains embeddings for all layers and $\mathbf{E}_{l}\in\Re^{N\times d_{e}}$, where $N$ is the number of experts and $d_{e}$ is the expert embedding dimension. The purpose of having the hypernetwork to generate $\mathbf{E}_{\text{all}}$ is to introduce inter-layer dependencies across different layers and operations. This design has been shown to accelerate the learning process in practice (Gao et al., 2024). More details are given in the Appendix A.
@@ 3.2 Expert Construction
In this section, we will talk about how to construct experts from MLP layers. In a decoder layer, the formulation of MLP is: $f_{\text{MLP}}(\mathbf{X})=\sigma(\mathbf{X}\mathbf{W}_{G})\odot(\mathbf{X}\mathbf{W}_{U})\mathbf{W}_{D}$, where matrices $\mathbf{W}_{U}\in\Re^{d\times d_{\text{mid}}}$, $\mathbf{W}_{G}\in\Re^{d\times d_{\text{mid}}}$ and $\mathbf{W}_{D}\in\Re^{d_{\text{mid}}\times d}$ denote up, gated, down projection matrices, $\sigma$ denotes nonlinear activation functions and $\odot$ denotes the Hadamard product (element-wise product).
Assume we want to have $N$ experts, under the setting of structural pruning, each expert can be represented by:
${}f_{\text{MLP}}^{i}(\mathbf{X})=\sigma(\mathbf{X}\mathbf{W}_{G}\mathbf{S}_{i})\odot(\mathbf{X}\mathbf{W}_{U}\mathbf{S}_{i})\mathbf{S}_{i}^{\top}\mathbf{W}_{D}$ (2)
where $i=1,\ldots,N$, $\mathbf{S}_{i}=\text{Diag}(\mathbf{s}_{i})$ ($\mathbf{s}_{i}\in\Re^{d_{\text{mid}}}$, $\mathbf{S}_{i}\in\Re^{d_{\text{mid}}\times d_{\text{mid}}}$), is a diagonal matrix with 0 or 1 and it is used to select a subset of weight vectors for the $i$th expert. After we have the formulation for each expert, we need to learn the configurations of each expert as shown below:
${}\mathbf{s}={\text{ST-GS}}(\text{Proj}_{\text{D}}^{\text{\tiny MLP}}(\mathbf{G}\mathbf{E})),\ \mathbf{G}={\text{ST-G}}(\text{Router}(\mathbf{X}))$ (3)
where $\mathbf{G}\in\Re^{T\times N}$ is the output of the router module, $\mathbf{E}$ ($l$ is omitted for clarity) is the expert embeddings, $\text{Proj}_{\text{D}}^{\text{\tiny MLP}}:\Re^{d_{e}}\to\Re^{d_{\text{mid}}}$ is a projection module to project the latent embedding to the MLP middle dimension, and ST-GS and ST-G are Straight-Through Gumbel-Sigmoid and Gumbel-Softmax functions respectively (Jang et al., 2016). Under this setting, $\mathbf{s}_{i}$ will contain retained positions (represented by $1$) for the $i$th expert and $\mathbf{G}$ contains one-hot routing decisions for different tokens in $\mathbf{X}$.
@@ 3.3 MHA top-K Routing
A MHA layer can be represented as: $f_{\text{MHA}}(\mathbf{X})=\sum_{i=1}^{H}\sigma_{s}\left({e}(\mathbf{X}\mathbf{W}_{Q,i}){e}^{\top}(\mathbf{X}\mathbf{W}_{K,i})\right)\mathbf{X}\mathbf{W}_{V,i}{\mathbf{W}_{O,i}}$, where $\mathbf{W}_{Q,i},\mathbf{W}_{K,i},\mathbf{W}_{V,i}\in\Re^{d\times\frac{d}{H}},\mathbf{W}_{O,i}\in\Re^{\frac{d}{H}\times d}$ are the query, key, value, and output matrices for each attention head, and $\mathbf{X}\in\Re^{T\times d}$ is the input hidden states. ${e}$ and $\sigma$ denote positional embedding and the softmax function.
@@
For MHA layers, we perform two kinds of pruning: dynamic top-K pruning and static pruning, both along the head dimension. More specifically, we can also insert selection matrices like MLP layers:
${}\begin{split}f_{\text{MHA}}(\mathbf{X}_{t})&=\sum_{i=1}^{H}[\sigma_{s}\left({e}(\mathbf{X}_{t}\mathbf{W}_{Q,i})\mathbf{S}_{0}\mathbf{S}_{0}^{\top}{e}^{\top}(\mathbf{X}_{t}\mathbf{W}_{K,i})\right)\\ &\mathbf{X}_{t}\mathbf{W}_{V,i}\mathbf{S}_{t}]\mathbf{S}_{t}{\mathbf{W}_{O,i}}\end{split}\vspace{-5pt}$ (4)
@@
where $\mathbf{S}_{0}\in\Re^{\frac{d}{H}\times\frac{d}{H}}$ is the shared selection matrix for static pruning of query and key matrices and $\mathbf{S}_{t}\in\Re^{\frac{d}{H}\times\frac{d}{H}}$ is the selection matrix for the value and output matrices of the $t$th token. We apply the same selection matrix across all heads, ensuring that all heads have the same head dimensions at the inference time. To generate the selection matrix, we calculate its diagonal vector $\mathbf{s}_{t}$ as:
${}\mathbf{s}_{t}={\text{ST-GS}}(\text{Proj}_{\text{D}}^{\text{\tiny MHA}}(\text{Proj}_{\text{E}}^{\text{\tiny MHA}}(\mathbf{X}_{t})+\frac{1}{N}\mathbf{1}^{\top}\mathbf{E})),$ (5)
@@
where $\mathbf{1}\in\Re^{N\times 1}$ is an all one vector, $\frac{1}{N}\mathbf{1}^{\top}\mathbf{E}$ represents the average expert embedding of size $d_{e}$, $\text{Proj}_{\text{D}}^{\text{\tiny MHA}}:\Re^{d_{e}}\to\Re^{\frac{d}{H}}$ is a project module to project the latent embedding to the head dimension and $\text{Proj}_{\text{E}}^{\text{\tiny MHA}}:\Re^{d}\to\Re^{d_{e}}$ is also a project module to project input tokens to the space of expert embeddings, and ST-GS is defined in Sec. 3.2. When $t=0$, we let $\mathbf{X}=0$, and $\mathbf{s}_{0}={\text{ST-GS}}(\text{Proj}_{\text{D}}^{\text{\tiny MHA}}(\frac{1}{N}\mathbf{1}^{\top}\mathbf{E}))$ since it is input independent.
@@
During training, the number of ones in $\mathbf{s}$ can vary freely. After training is complete, we compute $K=\text{round}(\frac{1}{T}\sum_{t=1}^{T}\sum_{i=1}^{\frac{d}{H}}\mathbf{s}_{t,i})$ for a subset of tokens and use it for top-K routing during inference. Note that the $k$ in top-k for $\mathbf{s}_{0}$ and $\mathbf{s}_{t}$ ($t\geq 1$) can be different, allowing for larger flexibility. Also, note that $\mathbf{s}_{0}$ needs to have specific structures to comply with the position embedding $e(\cdot)$, and more details are given in the Appendix A.2.

### v2

@@ Abstract
Large Language Models (LLMs) demonstrate remarkable capabilities but face deployment challenges due to their high computational demands. Traditional pruning methods reduce these costs by permanently removing parameters, which inevitably leads to performance degradation. To mitigate this issue, we propose ToMoE, a method that transforms dense LLMs into Mixture-of-Experts (MoE) models by uncovering experts inherently present within dense models, without requiring any weight updates. ToMoE leverages dynamic structural pruning to unify expert construction and router training in a single stage, achieving consistently strong performance. Remarkably, even without fine-tuning the model weights, ToMoE consistently outperforms state-of-the-art pruning and MoE techniques across Phi-2, LLaMA-2, LLaMA-3, and Qwen-2.5 models. The code for this paper is available at https://github.com/gaosh/ToMoE.
@@ 1 Introduction
Although LLMs demonstrate remarkable capacity to perform diverse tasks (brown2020language; kenton2019bert; raffel2020exploring; openai2022chatgpt; anthropic2023claude; dong2022survey; radford2019gpt2; kaplan2020scaling), their huge model size often limits their usability on devices with limited resources. As a result, considerable efforts (ma2023llm; ashkboos2023slicegpt; frantar2022gptq) are focused on minimizing the computational and memory costs of these models. Structural pruning (ma2023llm) has emerged as a promising solution to this challenge because, unlike unstructured pruning, it achieves compression without the need for specialized implementations. However, the problem with structural pruning methods is that they will substantially reduce the model capacity, resulting in an obvious performance gap compared to the dense model. The fine-tuning cost for even partially recovering this gap is tremendous.
To achieve a better trade-off between the number of parameters and performance, sparse Mixture of Experts (MoE) models (Shazeer2017; lepikhingshard) are designed to activate only a subset of the model’s parameters, corresponding to the selected experts. Recently proposed MoE models, such as DeepseekMoE (dai2024deepseekmoe), demonstrated that they can match the performance of dense models with a similar total parameter count while using a small number of active parameters. Following this lead, transforming dense models into MoE models could offer a promising approach to bridging the performance gap left by structural pruning methods. Unlike prior efforts to construct MoE models from dense models (zhang2022moefication; lee2024breaking; zhu2024llama), our findings reveal that MoE inherently exists within dense models and can be uncovered without updating model weights (continue pretraining). Specifically, we show that these experts can be identified through dynamic structural pruning. These results represent a novel contribution that has not been demonstrated in previous studies.
@@
The core idea of MoE models is conditional computation, where experts are dynamically selected based on input tokens. This concept aligns closely with dynamic pruning methods (gao2018dynamic), which make pruning decisions given input features. Leveraging this connection, we propose to construct MoE models from dense models by using dynamic structural pruning. Specifically, for Multi-Head self-Attention (MHA) layers, we apply top-K routing and static pruning for compression, while for MLP layers, we transform them into MoE layers using top-1 expert routing. The routing mechanism learned for dynamic structural pruning can be directly applied to serve as the routing module for MoE layers. With differentiable discrete operations, the MoE conversion process can be formulated as a differentiable dynamic pruning problem. With this formulation, we can efficiently convert a dense model to an MoE model at a cost similar to or lower than regular structural pruning methods. The comparison between our method, static pruning, and the original LLM is shown in Fig. 1.
Built upon the above findings and techniques, we proposed ToMoE to effectively convert dense LLMs to MoE models with dynamic pruning. The contributions of this work can be summarized as follows:
@@
  • Joint Optimization for Routing and Experts: The proposed method involves jointly optimizing routing modules and expert configurations by solving a regularized optimization problem. Our approach leverages differentiable operations to enable efficient and flexible MoE constructions.
  • Consistent Performance Improvements: Our method consistently outperforms state-of-the-art structural pruning and MoE construction techniques on various tasks while training only the router, without fine-tuning the model weights. This performance improvement is demonstrated across widely used public models such as Phi-2, LLaMA-2, LLaMA-3, and Qwen-2.5.
  • Detailed Analysis: We extensively analyze the resulting model from ToMoE across multiple perspectives, including parameter allocation, router behavior, and the ablation of different design components. We hope these analyses provide valuable insights and guidance for future research in this area.
@@ 2 Related Works
Figure 1: (a): The original LLM uses all parameters to process all the input text. (b): The static pruning for LLMs permanently removes model parameters, and the rest of the parameters are used to process all the input text. Our approach (c): LLMs with dynamic pruning use different sub-networks (illustrated by different colors) to process different tokens. We incorporate MoE to achieve a fixed expected budget in inference.
@@
Pruning: Structural pruning (li2017pruning; bert-surgeon; ma2023llm) is an attractive technique for real-world deployments since it removes redundant parameters to reduce model size without requiring specialized implementations. Structural pruning methods fall into two main categories: static pruning (cnn-static-pruning; molchanov2019importance; fang2023depgraph) and dynamic pruning (gao2018dynamic; chen2020dynamic; dynamic-context-pruning; kv-cache-compression). Static pruning removes parameters based on input-agnostic importance metrics. For example, LLM-Pruner (ma2023llm) eliminates non-essential coupled structures using gradient-based criteria. The problem with structural pruning is that it often creates a noticeable performance gap relative to dense models (ma2023llm; ashkboos2023slicegpt). In contrast, dynamic pruning removes weights based on input-dependent metrics. Early attempts for dynamic pruning (gao2018dynamic; chen2020dynamic) focus on Convolutional Neural Networks, where channels are selectively activated for input samples. Recent works, such as D-LLM (d-llm), incorporate the concept of conditional computation into LLMs by selectively skipping layers based on input tokens. The problem with dynamic pruning methods is that they do not have a fixed budget given different inputs, which creates problems when serving LLMs in a mini-batch setting or in the prefilling stage. Our method, on the other hand, converts the dense LLM to a sparse MoE model with a fixed per-token budget.
Another line of research applies contextual sparsity for LLMs (pmlr-v202-liu23am; zheng2024learn; lee2024cats), where neurons are selectively activated given certain conditions. Although there are some promising results, they are generally more difficult to achieve better inference efficiency due to their irregular sparsity patterns. In contrast, MoE models have more comprehensive support from the system side, making them a popular choice for scaling up the model. Thus, our method mainly focuses on converting dense models to MoE models.
@@
MoE: Sparse Mixture-of-Experts (MoE) models improve upon pure structural pruning by maintaining or even enhancing model capacity without a proportional increase in computational cost. For instance, Sparsely-Gated MoE (Shazeer2017) employs a trainable gating network to select a small subset of experts for each input, enabling the model to scale to thousands of experts efficiently (lepikhingshard). More recent methods like DeepSeekMoE (dai2024deepseekmoe) further address expert specialization, matching dense-model performance with a similar number of activated parameters. Previous methods constructing MoE from the dense model (zhang2022moefication; lee2024breaking; zhu2024llama) separate the expert construction and router training into two distinct stages, often leading to sub-optimal performance. In contrast, our method integrates expert construction directly into the pruning process, treating it as a unified step with router learning and thereby largely improving the performance without fine-tuning.
@@ 3 ToMoE
Most recent LLMs, like GPT (radford2018gpt), LLaMA (touvron2023llama), etc., adapt decoder-only architectures and thus our method focuses on decoder-only architectures. A typical decoder block consists of Multi-Head Attention (MHA) and Multi-Layer Perceptron (MLP) layers. For clarity, we denote the sequence length by $T$, the hidden dimension by $d$, the MLP intermediate dimension by $d_{\text{mid}}$, and the number of attention heads by $H$.
To reduce the computational costs of the decoder-only architecture, we propose to convert the original model into MoE models. For MHA layers, we utilize top-K routing and static pruning along the head dimension $\frac{d}{H}$. Top-K routing and static pruning for MHA layers ensure that, during prefilling or model serving, all tokens maintain the same head dimension, enabling parallel processing. For MLP layers, our approach transforms them into MoE layers along the MLP middle dimension $d_{\text{mid}}$ and employs top-1 routing. A key distinction between our method and previous dynamic pruning approaches is that the converted model maintains consistent computational costs for all inputs. This property could be crucial for efficient processing.
@@ 3.1 Expert Embeddings
Inspired by the recent success of using hypernetworks (ha2016hypernetworks; ganjdanesh2024not; gaodisp) to generate pruning decisions, we adopt a hypernetwork to generate expert embeddings:
$\mathbf{E}_{\text{all}}=\text{HN(z)},$ (1)
where $z$ is the input to the hypernetwork drawn from a random distribution, and $\mathbf{E}_{\text{all}}=[\mathbf{E}_{1},\cdots,\mathbf{E}_{l},\cdots,\mathbf{E}_{L}]$ contains embeddings for all layers and $\mathbf{E}_{l}\in\mathbb{R}^{N\times d_{e}}$, where $N$ is the number of experts and $d_{e}$ is the expert embedding dimension. Each embedding $\mathbf{E}_{l,i}$ will then be used to generate the configurations of experts. The purpose of having the hypernetwork to generate $\mathbf{E}_{\text{all}}$ is to introduce inter-layer dependencies across different layers and operations. This design has been shown to accelerate the learning process in practice (gaodisp). More details are given in the Appendix A.
@@ 3.2 Expert Construction
In this section, we will talk about how to construct experts from MLP layers. In a decoder layer, the formulation of MLP is: $f_{\text{MLP}}(\mathbf{X})=\sigma(\mathbf{X}\mathbf{W}_{G})\odot(\mathbf{X}\mathbf{W}_{U})\mathbf{W}_{D}$, where matrices $\mathbf{W}_{U}\in\mathbb{R}^{d\times d_{\text{mid}}}$, $\mathbf{W}_{G}\in\mathbb{R}^{d\times d_{\text{mid}}}$ and $\mathbf{W}_{D}\in\mathbb{R}^{d_{\text{mid}}\times d}$ denote up, gated, and down projection matrices. In addition, $\sigma$ denotes nonlinear activation functions and $\odot$ denotes the Hadamard product (element-wise product).
Assume the target is to use $N$ experts, under the setting of structural pruning, each expert can be represented by:
$f_{\text{MLP}}^{i}(\mathbf{X}_{t})=\sigma(\mathbf{X}_{t}\mathbf{W}_{G}\mathbf{S}_{i})\odot(\mathbf{X}_{t}\mathbf{W}_{U}\mathbf{S}_{i})\mathbf{S}_{i}^{\top}\mathbf{W}_{D},$ (2)
where $i=1,\cdots,N$, and $\mathbf{S}_{i}=\text{Diag}(\mathbf{s}_{i})$ ($\mathbf{s}_{i}\in\mathbb{R}^{d_{\text{mid}}}$, $\mathbf{S}_{i}\in\mathbb{R}^{d_{\text{mid}}\times d_{\text{mid}}}$), is a binary diagonal matrix selecting a subset of weight vectors for the $i$th expert. $\mathbf{X}_{t}$ is the $t$th token, which is assumed to be routed to the $i$th expert. Once each expert is formulated, its configuration is learned as follows:
$\mathbf{s}={\text{ST‑GSig}}(\text{Proj}_{\text{D}}^{\text{\tiny MLP}}(\mathbf{G}\mathbf{E})),\ \mathbf{G}={\text{ST‑GSmax}}(\text{Router}(\mathbf{X})),$ (3)
where $\mathbf{G}\in\mathbb{R}^{T\times N}$ is the output of the router module, $\mathbf{E}$ ($l$ is omitted for clarity) is the expert embeddings, $\text{Proj}_{\text{D}}^{\text{\tiny MLP}}:\mathbb{R}^{d_{e}}\to\mathbb{R}^{d_{\text{mid}}}$ is a projection module to project the latent embedding to the MLP middle dimension, $\text{Router}(\cdot):\mathbb{R}^{d_{e}}\to\mathbb{R}^{N}$ is the router module that maps the inputs to an $N$-dimensional routing score vector for expert selection, and ST‑GSig and ST-GSmax are Straight-Through Gumbel-Sigmoid and Gumbel-Softmax functions respectively (jang2016categorical). Under this setting, $\mathbf{s}_{i}$ will contain retained positions (represented by $1$) for the $i$th expert, and $\mathbf{G}$ contains one-hot routing decisions for tokens in $\mathbf{X}$.
@@ 3.3 MHA top-K Routing
An MHA layer can be represented as $f_{\text{MHA}}(\mathbf{X})=\sum_{i=1}^{H}\sigma_{s}\left(e(\mathbf{X}\mathbf{W}_{Q,i})e^{\top}(\mathbf{X}\mathbf{W}_{K,i})\right)\mathbf{X}\mathbf{W}_{V,i}\mathbf{W}_{O,i}$, where $\mathbf{W}_{Q,i},\mathbf{W}_{K,i},\mathbf{W}_{V,i}\in\mathbb{R}^{d\times\frac{d}{H}}$, $\mathbf{W}_{O,i}\in\mathbb{R}^{\frac{d}{H}\times d}$ are the query, key, value, and output matrices for each attention head, and $\mathbf{X}\in\mathbb{R}^{T\times d}$ is the input hidden states. e{e} and $\sigma_{s}$ denote positional embedding and the softmax function.
@@
For MHA layers, we perform two kinds of pruning: dynamic top-K pruning and static pruning, both along the head dimension. Justifications regarding the design choice are provided in the Appendix F. Like MLP layers, we also insert selection matrices:
$f_{\text{MHA}}(\mathbf{X}_{t})=\sum_{i=1}^{H}\left[\sigma_{s}\left(e(\mathbf{X}_{t}\mathbf{W}_{Q,i})\mathbf{S}_{0}\mathbf{S}_{0}^{\top}e^{\top}(\mathbf{X}_{t}\mathbf{W}_{K,i})\right)\mathbf{X}_{t}\mathbf{W}_{V,i}\mathbf{S}_{t}\right]\mathbf{S}_{t}^{\top}\mathbf{W}_{O,i},$ (4)
@@
where $\mathbf{S}_{0},\mathbf{S}_{t}\in\mathbb{R}^{\frac{d}{H}\times\frac{d}{H}}$ are selection matrices. $\mathbf{S}_{0}$ is the shared selection matrix for static pruning of query and key matrices, while $\mathbf{S}_{t}$ is the token-specific selection matrix for the value and output matrices of the $t$-th token. We apply the same selection matrix across all heads, ensuring that all heads have the same head dimensions at inference time. To generate the selection matrix, we calculate its diagonal vector $\mathbf{s}_{t}$ as:
$\mathbf{s}_{t}={\text{ST‑GSig}}(\text{Proj}_{\text{D}}^{\text{\tiny MHA}}(\text{Proj}_{\text{E}}^{\text{\tiny MHA}}(\mathbf{X}_{t})+\frac{1}{N}\mathbf{1}^{\top}\mathbf{E})),$ (5)
@@
where $\mathbf{1}\in\mathbb{R}^{N}$ is an all-one vector, $\frac{1}{N}\mathbf{1}^{\top}\mathbf{E}$ represents the average expert embedding of size $d_{e}$, $\text{Proj}_{\text{D}}^{\text{\tiny MHA}}:\mathbb{R}^{d_{e}}\to\mathbb{R}^{\frac{d}{H}}$ is a projection module to map the latent embedding to the head dimension and $\text{Proj}_{\text{E}}^{\text{\tiny MHA}}:\mathbb{R}^{d}\to\mathbb{R}^{d_{e}}$ is also a projection module to project input tokens to the space of expert embeddings, and ST‑GSig is defined in Sec. 3.2. When $t=0$, we initialize $\mathbf{X}=0$, and set $\mathbf{s}_{0}={\text{ST‑GSig}}(\text{Proj}_{\text{D}}^{\text{\tiny MHA}}(\frac{1}{N}\mathbf{1}^{\top}\mathbf{E}))$, since it is input independent.
@@
During training, the number of ones in $\mathbf{s}$ can vary freely. After training is complete, we compute $K=\text{round}(\frac{1}{T}\sum_{t=1}^{T}\sum_{i=1}^{\frac{d}{H}}\mathbf{s}_{t,i})$ for a subset of tokens and use it for top-K routing during inference. Note that the $k$ in top-K for $\mathbf{s}_{0}$ and $\mathbf{s}_{t}$ ($t\geq 1$) can be different, allowing for larger flexibility. Also, note that $\mathbf{s}_{0}$ must follow specific structural constraints to be compatible with the position embedding $e(\cdot)$, and more details can be found in Appendix A.2.
3.4 Regularizations for MoE Constructions
In Sec. 3.2 and Sec. 3.3, we briefly introduced the design space for constructing MoE models using dynamic structural pruning. In this subsection, we will introduce regularizations customized to the characteristics of MoE models.
where $\bigcup$ is the union operator, and $\mathbf{u}$ is the union of all kept positions for each token. For MLP layers, it can be calculated similarly. We then push $\frac{\sum\mathbf{u}}{|\mathbf{u}|}$ ($|\mathbf{u}|$ represents the size of $\mathbf{u}$) to 1:
$\mathcal{R}_{\text{U}}=\frac{1}{L}\sum_{l=1}^{L}f_{\text{reg}}(\frac{\sum\mathbf{u}_{l}}{|\mathbf{u}_{l}|},1),$ (7)
where $f_{\text{reg}}(\cdot,\cdot)$ can be any regression loss functions, and we will choose $f_{\text{reg}}$ later.
Parameter Regularization. For a sparse MoE model, we also need to control the number of active parameters given the provided budget. To achieve this goal, we can directly penalize the maximum width across different experts. We choose the maximum width over experts instead of the mean, median, or other alternatives because the maximum provides precise control over the upper bound of the number of active parameters.
Denote the width of a layer as $d_{l}^{*}$, where $\in\{\text{MLP},\text{MHA}\}$. For MLP layers, it can be calculated by $d_{l}^{\text{ MLP}}=\max(\mathbf{s}\mathbf{1}_{d_{\text{mid}}})$, where $\mathbf{1}_{d_{\text{mid}}}\in\mathbb{R}^{d_{\text{mid}}}$ is an all-one vector of size $d_{\text{mid}}$. $\mathbf{s}\mathbf{1}_{\text{mid}}$ produces the width of all experts, and $d_{l}^{\text{ MLP}}$ represents the maximum width across all experts. The width of MHA layers can be calculated similarly. Based on $d_{l}^{*}$, we can calculate the number of active parameters in the model $\text{T}(\mathbf{d_{\text{MoE}}})$, where $\mathbf{d_{\text{MoE}}}=[d_{1}^{*},\cdots,d_{L}^{*}]$. To push the number of active parameters to a predefined rate $p$, the following objective is applied:
$\mathcal{R}_{\text{P}}=f_{\text{reg}}(\text{T}(\mathbf{d_{\text{MoE}}}),p\text{T}_{\text{total}}),$ (8)
where $\text{T}_{\text{total}}$ is the total number of parameters, and $p\in(0,1]$ represents the ratios of the active parameters. For $f_{\text{reg}}$ in Eq. 8 and Eq. 7, the following function $f_{\text{reg}}$ is used:
$f_{\text{reg}}(x,y)=\log(\max(x,y)/\min(x,y)).$
Load Balancing Regularization. When determining the configurations of experts, we also apply the load balancing regularization to encourage a balanced load across experts (lepikhin2021gshard; fedus2022switch). The load balancing loss from the Switch Transformer (fedus2022switch) is adopted:
$\mathcal{R}_{\text{L}}=N\sum_{i=1}^{N}F_{i}P_{i},$ (9)
where $F_{i}=\frac{1}{T}\sum_{t=1}^{T}\mathbbm{1}(\mathbf{G}_{t,i}=1)$. The indicator function $\mathbbm{1}(\cdot)$ returns $1$ if the condition is true and 0 otherwise. $F_{i}$ represents the fraction of tokens assigned to the $i$-th expert. $P_{i}=\frac{1}{T}\sum\text{GSmax}({G}_{t,i})$, where ${G}=\text{Router}(\mathbf{X})$ is the router output before ST‑GSmax. $P_{i}$ is the fraction of the router probability allocated for the $i$-th expert. $\mathcal{R}_{\text{L}}$ will encourages uniform routing across different experts as shown in (fedus2022switch). The combination of Eq. 8 and Eq. 9 creates an interesting phenomenon where they encourage uniform allocation of width among experts.